Consider the following boundary-value problem

$$ y^{\prime \prime}-2 y^{\prime}+y=x^{2}-1, y(0)=2, \quad y(1)=4 $$

Apply the linear shooting method and the Euler method with step size of \(\frac{1}{3}\) to marks) approximate the solution of the problem.
Answer #1

SOLUTION:

☺ (0) Solution: - Given thot 4-24}+y= 222 4109=2; y(i)=4. bet ; y(0)=2 VP y=p= dig op taf,(2,4,7) de = y= 20-4y+z2_1 ; Géo)246 - 13.926 = 4, 27 Assume yro) =4 and neglect of again let us assume global Now î পু। yi Pi 1 0 2 0 r Y 5/3 2 2/3 513 -23/3 - 3 m-4- (4-13.926) 12.544 ma 1.626